OpenStudy (anonymous):

PLEASE HELP!! A local florist charges $8 for a red rose and $10 for a white rose. He sold twice as many red roses as white roses for a total income of $8,554 for the day. How many roses did he sell all together?

OpenStudy (anonymous):

its either a. 329 b.658 c.987 d.8554

OpenStudy (anonymous):

x=8554/26

OpenStudy (anonymous):

how you get 26

OpenStudy (anonymous):

8*2*x+10*x=8554

OpenStudy (anonymous):

you need to solve the equation

OpenStudy (anonymous):

16x+10*x=8554

OpenStudy (anonymous):

26x=8554 which is 329

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you're welcome